OpenStudy (anonymous):

Matthew purchased a washer and dryer for $1,965 using an 18-month deferred payment plan with an interest rate of 18.43%. What is the balance after the deferment period if payments of $65 were made each month? $1,965.00 $2,585.30 $795.00 $1,415.30

OpenStudy (amistre64):

this is just a balance equation ...

OpenStudy (amistre64):

1965k^(18) - 65(1-k^(18))/(1-k)

OpenStudy (amistre64):

instead of having to pay a set minimum payment to zero out the loan in a certain time span; they just created their own payment structure

OpenStudy (dumbcow):

@amistre64 , it doesn't match any of the answers though i worked it out twice and got 1249

OpenStudy (amistre64):

i get 1249 as well

OpenStudy (dumbcow):

does deferred mean no interest is applied for those 18 months the 65 is just simply deducted from balance?

OpenStudy (amistre64):

not unless it says no interest accruing; it means that there is no payments necessary during the period

OpenStudy (dumbcow):

yeah and thats never the case...never do deferred loans haha

OpenStudy (amistre64):

Bo = Bo B1 = Bo(k) - P B2 = (Bo(k) - P)k - P = Bok^2 -Pk -P B3 = Bok^3 -Pk^2 -Pk -P Bn = Bok^n -Pk^(n-1) -Pk^2 -Pk -P \[B_n=B_ok^n-P\frac{1-k^n}{1-k}\] allowing for k to clean up that compounding mess

OpenStudy (amistre64):

since n is measured in months; n = 18 in this case

OpenStudy (dumbcow):

based on the choices it must be option D none of others make sense...balance would change and be less than if no payments were made but greater than option where no interest accrues

OpenStudy (amistre64):

i agree

OpenStudy (amistre64):

then that would be a defered period of about 14 to 15 months :)